# Direct imaging and Talbot Phase Imaging at the Matter in Extreme Conditions X-ray Imager of LCLS

### v2

@@ Abstract
The last decade has shown the great potential that X-ray Free Electron Lasers have to study High Energy Density matter. Experiments at FELs have made significant breakthroughs in Shock Physics and Dynamic Diffraction, Dense Plasma Physics and Warm Dense Matter Science, using techniques such as isochoric heating, inelastic scattering, small angle scattering and diffraction. In addition, and complementary to these techniques, the coherent properties of the beam can be used to image HED samples with high precision. We present new imaging diagnostics and techniques developed at the MEC instrument at LCLS over the last few years. We briefly show result of a previously used Phase Contrast Imaging setup, where the X-ray beam propagates from the target to a camera some distance away revealing its phase, as well as from a new direct imaging approach where the target is re-imaged on the camera with 300nm resolution. Last, we show a new Talbot Phase Imaging method that allows for both measuring the phase and intensity changes of the x-rays introduced by a target with sub-micron resolution.
@@ I Introduction
HED science has in the past made ample use of X-ray radiography and imaging to investigate a variety of phenomena, such as the implosions of Inertial Confinement Fusion capsules Katayama et al. (1993); Kalantar et al. (1997); Marshall et al. (2009); Hicks et al. (2010), X-pinch plasmas Shelkovenko et al. (2001) and the general hydrodynamics of evolution of HED targets. Landen et al. (2001). Since the dynamics of these phenomena evolve on very fast time scales (from fs to ns), short X-ray pulses are needed to obtain useful data. Historically, fast X-ray sources generated by the interaction of high power laser with backlighter targets have been employed, with typical spatial and temporal resolution on the order of tens of micrometers and hundreds of picoseconds or more, respectively. Techniques to use Phase Contrast Imaging with backlighters have also been developed Montgomery et al. (2004); Kozioziemski et al. (2005); Koch et al. (2009); Ping et al. (2011), albeit with limited resolutions compared to synchrotrons, where sub-micron resolutions are standard Nugent et al. (1996); Cloetens et al. (1999).
The start of operation of LCLS Emma et al. (2010) in 2009 opened new possibilities for X-ray imaging in HED science. The coherence properties of the beam allow for imaging with sub-micron spatial resolution, while the short, bright pulses allow temporal resolution of tens of femtoseconds. These combined properties allow for single-shot X-ray imaging techniques to emerge as a leading technique to further our understanding of complex HED phenomena. At LCLS, the Matter in Extreme Conditions end-station Nagler et al. (2015) is specifically tailored to field experiments in HED science and to foster this field of research, the different imaging techniques and instrumens have been developed, and are now available to its user community.
In this paper, we describe the use and performances of the MEC Xray Imager instrument (MXI) and the different diagnostic imaging techniques that make use of it, then show some preliminary proof of principle experiments to showcase its potential for HED science.
@@ II Instrument overview
The MXI instrument is designed for use in the MEC vacuum chamber, details of which can be found in Nagler et al. (2015). It uses compound refractive Beryllium lenses to focus the X-ray beam to sub-micron spot sizes. The capabilities of such lens stack have been describe extensively in the literature Snigirev et al. (1996); Lengeler et al. (1999a, b, 2002), and have been used for this purpose at MEC before Nagler et al. (2016); Schropp et al. (2013); Seiboth et al. (2017). A stack of lenses is placed on alignment stages that position the lenses in the beam. An 3D model is shown in Figure 1. Up to three stacks of lenses can be placed on the instrument simultaneously, to allow using different photon energies and/or focal lengths during an experiment without venting the MEC vacuum chamber. The front stack can accommodate up to 100 lenses, necessary for the higher photon energies, while the other two stacks are limited to 40 lenses. The lens stacks are mounted on a hexapod, allowing for precise alignment with six degrees of freedom, and a translation stage of 500 mm to change the distance between the lens stack and the object plane. Since stacks of Be lenses are known to have aberrations, which can limit their use in imaging and phase contrast application, a phase plate that corrects these aberrationsSeiboth et al. (2017) can be placed 10 mm in front of the lens set. The phase plate can be aligned in the horizontal and vertical direction with two small linear stages, which can also completely retract the corrector. The MXI can be placed either upstream of the sample (see Sec.III) for phase contrast imaging or downstream (see Sec.IV) for direct imaging of a sample.
@@
In a series of experiments used to date, samples are driven by the MEC short or long pulse laser orthogonally to the imaging direction (e.g. a shock is imaged traveling through a sample). To this end a sample holder allowing such geometry is available (see Fig. 2).
@@
The samples are aligned using two long distance microscopes (see Nagler et al. (2015) for details), which view the sample at orthogonal directions, and allow for precise and repeatable alignment. The samples are mounted on pillar that are screwed to the sample holder. For use in shock experiment, specialty pillars that allow for VISAR Barker and Hollenbach (1972) measurements on thin samples are available.
@@
The image is recorded on an X-ray camera that sits behind the MEC chamber. In principle, the camera can be placed between 1.3m to 4.3m behind the sample plane. In general, the farthest distance is preferred to have the largest magnification, although some application may benefit from the shorter distance (i.e. ptychography Schropp et al. (2010)). For most experiments, we have used an Optique Peter X-ray microscope, consisting of a Ce:YAG scintillator and optical microscope, coupled to a Andor Zyla optical camera, which can have a resolution of 2 m. Direct imaging cameras (e.g. Epix or Jungfrau) can be used instead, however the resolution is limited by the pixel size (typically 50m for Epix10k and 75um for Jungfrau). A framing camera, capable of recording up to four images spaced as little as 350ps apart can also be used. An advantage of direct detection compared to scintillator based detection is that the camera’s are orders of magnitude more sensitive, and therefore can be used with thicker, more absorbent samples.
@@ III Phase Contrast Imaging geometry
Here, the LCLS beam is focused down with the MXI inside the MEC vacuum chamber with a lens set of focal length , to a spot size of typically 100 nm. A certain distance, x, behind the focus, a sample is placed in the divergent beam. The x-ray beam eventually hits a detector that sits a distance L behind the sample, leading to a geometric magnification of the sample of L/x on the detector. Due to the long propagation distance of the X-ray, phase difference in the sample can readily be seen in the image (e.g. clear fringes appear at sharp phase boundaries) For this reason this geometry is called the Phase Constrast Imaging (PCI) geometry at MEC. Iterative algorithms can be used to retrieve the phase, and if this is successful the full complex transmission function (i.e. absorption and phase shift) can be calculated. Experiment in this geometries have been fielded at MEC before yielding scientific results on shockwaves is diamond Schropp et al. (2015) and the kinetics of phase transformations in silicon Brown et al. (2019). The setup can be easily combined with X-ray diffraction. Since moving the lenses of the MXI closer to the target reduced the field of view of the image, and therefore the area where X-rays hit the target, diffraction from selected area of only microns in size can be obtained to determine different phases in the image, as was done in Brown et al. (2019). A typical image of a laser-generated shock wave travelling through a sample can be seen in Fig. 4 and many more examples in have been published in Schropp et al. (2015); Nagler et al. (2015, 2016)
@@ IV Direct Imaging geometry
While the PCI geometry described above has great potential when a full phase retrieval can be performed, this has been in most cases elusive at MEC. Since without such a phase retrieval, the resolution can be severely limited, an alternative setup that directly images a sample on the detector has been developed and used at MEC. A sketch of the setup is shown in Figure 5
@@
In this geometry the Be lenses are placed behind the sample such that the distance to the sample,, and the distance to the camera, , match the well know lens formula , with the focal length of the Be lenses. Such a setup allows for imaging intensity contrast of the sample with sub-micron resolution, even when phase reconstruction is not possible. Result showing the resolution better than 250nm that was attained by this setup are shown in Figure 6.
While in theory, a perfect image of the intensity at the sample location (and therefore the absorption of the sample) is obtained, this is not the case in practise since the imaging is not perfect. First and foremost the image of the x-rays on the detector when no sample is present has significant structure that are introduced by the Be lens stack. Some of this can be treated by a simple white field correction (i.e. dividing the image of the sample by the white field of image when no sample is present) or by more soffisticated analysis method such as Principal Component Analysis. This is the analysis that is performed in the images in Figures 4, 6 and 7. However, this method does not take into account the phase introduced by the structure in the be lenses, and therefore doesn’t work very well in areas of he image where strong phase changes are expected (e.g. at a shock front). This effect gets worse when lenses upstream of the MEC target chamber are placed into the beam to pre-focus the x-rays onto the part of the sample that is imaged, to increase signal level. Furthermore, both wavefront aberrations and chromatic aberrations are present in the imaging lens stack. The chromatic aberration can be mitigated by a monochromator, however this is not always feasible: the loss of photons due to the monochromator can reduce the number of photon below what is necessary to get an adequate signal on the camera. The wavefront aberration can be reduced with the appropriate phase-correcter if it is available; however this correction is never perfect. Furthermore, the limited aperture of the lenses give an intrinsic limit to the imaging that can be achieved. Since the LCLS light source is virtually fully spatially coherent, these imaging imperfection tend to show up as fringes, especially around abrupt changes in density where jumps in the phase of the x-rays are created (e.g. sample edges, edges of wires, shock front, etc.) Furthermore, samples can be thicker than the theoretical Rayleigh length of the focus (or depth of focus of the imaging), leading to significant propagation effects inside the sample. The commonly used assumption that the sample can be modelled by an infinitely thin absorption and phase masks breaks down. These effects can all present significant challenges to the interpretation of and modelling of the data. A full treatment of these analysis challenges is outside the scope of this paper.
@@
The MXI instrument was tested in the direct imaging geometry at 8.2 keV and 18 keV photon energy. For 8.2 keV, a set of 25 Beryllium lenses with a curvature of 50 m were used. The set has a focal length of 204 mm and was placed 214 mm after the target, leading to an image at the detector, placed 4.1 m behind the lenses, with a magnification of approximately 20. The position of the lenses was tuned to get the best imaging resolution, using resolution targets. Images of the resolution target are shown in Fig.6. Line widths as small as 250nm can be clearly resolved.
@@
At photon energies as high as 18keV, the setup and alignment is complicated by the fact that many more Be lenses need to be used, making alignment harder, and increasing the lens aberrations. In addition, the shortest focal lengths that can be used are generally larger, leading to smaller magnification. Nevertheless, the higher photon energy allow the used of either thicker targets, or targets containing higher Z-elements, while maintaining enough transmission to collect good images. We tested the system with a lens stack of 98 lenses with a 50m curvature at their apex, leading to a focal length of 266mm and an magnification of 15. Measurement on resolution targets showed a resolution similar to 8.2keV. At 18keV, this is significantly bigger than the theoretical diffraction limit, due to a combination of both geometric and chromatic aberration of the Be lenses.
@@
Some examples of the use of this imaging geometry are shown below. In Figure 7 a shockwave in a silicon sample, driven by the MEC long pulse laser is depicted. Both an elastic precursor, and phase change to can be clearly distinguished.
@@
The imaging setup can also be used to measure the spotsize of the standard beryllium focusing lenses that are in the MEC beamline. These lenses sit approximately 4.2m upstream of the center of the MEC target chamber (see Nagler et al. (2015) for details), and therefore a stack with focal length between 4.0m and 4.4m is generally used for standard MEC experiment, with a best focus around 1.2 m. Wavefront measurement have in the past been performed of such lens stack, showing significant spherical aberations, and therefore ring-patterns around the focus Seaberg et al. (2019). The MXI direct imaging setup can however be used to directly image the intensity profile of that impinges on the sample that these lenses generate. In Figure 8 we show the intensity profile around the focus of our beamline lenses as imaged by the MXI instrument. The airy-like ring pattern can clearly be seen.
@@
V Talbot Phase Imaging (TPI)
Talbot shearing interferometry has been widely used at Free electron laser to determine the wavefront of the of the x-ray beam, and determine the its focal properties in a single shot Seaberg et al. (2019); Liu et al. (2018); Matsuyama et al. (2012). The method uses a two dimensional phase diffraction grating placed in the beam. At a specific distance behind the grating an ’image’ of the grating is reproduced, and analysis of the distortions, typically using Fourier methods Takeda et al. (1982) yields the differential phase in two orthogonal dimensions. These differential phases can be integrated to yield the full wavefront of the beam, and backpropagating to the focus yields the focal properties. It has been in important technique in quantifying the focus in experiments where x-ray foci of order 100nm were required.
We have have combined this Talbot phase measurement technique with the direct imaging. The setup, which can be seen in Figure 9, is essentially the same as the direct imaging setup, with the addition of a Talbot grating. A checkerboard -phase grating with a pitch of 22.5 m was placed at the first Talbot distance in front of the camera.
The raw image of the a shock propagating through Silicon with this setup is shown in Figure 10.a). Both the elastic precursor as well as the phase transformation to denser shocked state of Si can be seen in the raw image. Overlayed on this image is the grid pattern introduced by the Talbot grating (an zoom region is shown in the inset) that is used measure the phase. The 2-dimensional fringe pattern is analysed by standard Fourier methods Takeda et al. (1982). Since Talbot interferometry is a shearing interferometry method, the Fourier analysis results in the differential phase in the horizontal or vertical direction depending on which Fourier peak is chosen. Since at the first Talbot order the shear is small, we can approximate this as:
There are many ways to integrate the differential phase and get the phase . A fast an easy method Morgan et al. (2012) uses the Fourier relation:

### v4

# X-ray Microscopy and Talbot Imaging with the Matter in Extreme Conditions X-ray Imager at LCLS

@@ Abstract
The last decade has shown the great potential that X-ray Free Electron Lasers (FEL) have to study High Energy Density (HED) matter. Experiments at FELs have made significant breakthroughs in Shock Physics and Dynamic Diffraction, Dense Plasma Physics and Warm Dense Matter Science, using techniques such as isochoric heating, inelastic scattering, small angle scattering and x-ray diffraction. In addition, and complementary to these techniques, the coherent properties of the FEL beam can be used to image HED samples with high fidelity. We present new imaging diagnostics and techniques developed at the Matter in Extreme Conditions (MEC) instrument at Linac Coherent Light Source (LCLS) over the last few years. We show results in Phase Contrast Imaging geometry, where the X-ray beam propagates from the target to a camera revealing its phase, as well as in Direct Imaging geometry, where a real image of the sample plane is produced in the camera with a spatial resolution down to 200 nm. Last, we show an implementation of the Talbot Imaging method allowing both x-ray phase and intensity measurements change introduced by a target with sub-micron resolution.
@@ I Introduction
High Energy Density (HED) science has traditionally made ample use of X-ray radiography and imaging to investigate a variety of phenomena, such as the implosions of inertial confinement fusion (ICF) capsules Katayama et al. (1993); Kalantar et al. (1997); Marshall et al. (2009); Hicks et al. (2010); Dewald et al. (2018); Do et al. (2022), X-pinch plasmas Shelkovenko et al. (2001) and the general hydrodynamics of evolution of HED targets. Landen et al. (2001); Oliver et al. (2022). Since the dynamics of these phenomena evolve on very fast time scales (from femtoseconds to nanoseconds), ultrashort X-ray pulses are needed to obtain quantitative data. Historically, X-ray sources generated by the interaction of high power laser with backlighter targets have been employed, with typical spatial and temporal resolution on the order of tens of micrometers and hundreds of picoseconds or more Landen et al. (2001); Morace et al. (2014) Phase Contrast Imaging techniques with backlighters have also been developed Montgomery et al. (2004); Kozioziemski et al. (2005); Koch et al. (2009); Ping et al. (2011), albeit with limited resolutions compared to synchrotrons, where sub-micron resolutions are standard Nugent et al. (1996); Cloetens et al. (1999).
The start of operation of LCLS Emma et al. (2010) in 2009 opened new possibilities for X-ray imaging in HED science. The coherent properties of the beam allow for imaging with sub-micron spatial resolution, while the short, bright pulses allow temporal resolution of tens of femtoseconds. These combined properties allow for single-shot X-ray imaging techniques to emerge as a leading technique to further our understanding of complex HED phenomena. At LCLS, the Matter in Extreme Conditions end-station Nagler et al. (2015) is specifically tailored to field experiments in HED science. To foster this field of research, the different imaging techniques and instruments have been developed, and are available to its user community.
In this paper, we describe the use and performances of the MEC Xray Imager (MXI) and the different diagnostic imaging techniques that make use of it, and present preliminary proof of principle experiments to showcase its potential for HED science.
@@ II Instrument overview
The MXI instrument has two main components: an in-vacuum X-ray lens stack holder and an in-air optical microscope, as detailed in Figure 1.
@@
The in-vacuum component is designed for use in the MEC target chamber (technical details on the endstation can be found in Nagler et al. (2015)). It uses compound refractive Beryllium lenses (Be CRL) to focus the X-ray beam to sub-micron spot sizes. The capabilities of such lens stacks have been described extensively in the literature Snigirev et al. (1996); Lengeler et al. (1999a, b, 2002), and have been used for this purpose at MEC Nagler et al. (2016); Schropp et al. (2013); Seiboth et al. (2017). A stack of lenses is placed on alignment stages that positions the lenses in the beam. Up to three stacks of lenses can be placed on the instrument simultaneously, to allow usage of different photon energies and/or focal lengths during an experiment without venting the MEC vacuum chamber. The front stack can accommodate up to 100 lenses, necessary for the higher photon energies. The lens stacks are mounted on a hexapod, allowing for precise alignment with six degrees of freedom: the translations are aligned with the LCLS orthogonal coordinate system and rotation around each of these axis. The hexapod is attached to a linear translation stage with 500 mm travel range to change the distance between the lens stack and the object plane along the LCLS X-ray beam.
Two thin pins are precisely located with respect to the Be CRL stacks on the lens platform and are far from each other along the FEL direction to maximize alignment accuracy. They are initially used for the pre-alignment in air with the reference optical laser collinear with the FEL, and later, in vacuum conditions using the FEL itself via their shadowgraphy on the detector. These pins are critical to allow for a fast while precise alignment, typically achieved in less than thirty minutes.
Since stacks of Be lenses are known to have aberrations, which can limit their use in imaging and phase contrast application, a phase plate that corrects these aberrations Seiboth et al. (2017) can be placed 10 mm in front of the lens set. The phase plate is aligned in the horizontal and vertical direction with two small linear stages, which can also completely retract the corrector. The MXI lens sets can be placed either upstream of the sample (see Sec.III) for phase contrast imaging or downstream (see Sec.IV) for direct imaging of a sample.
@@
In typical HED experiments, samples are driven by the MEC short or long pulse laser orthogonally to the imaging direction (e.g., a shock is imaged traveling through a sample). To this end a sample holder allowing such geometry is available (see Fig. 2).
@@
The samples are aligned to TCC with better than 20 m accuracy using two long-distance telescopes (see Nagler et al. (2015) for details), which view the sample at orthogonal directions. The MXI can be used to further improve the alignment accuracy down to better than 1 m in the plan perpendicular to the FEL. The samples are mounted on a pillar that is screwed to the sample holder. The pillars are designed to accommodate for line of sight of additional scientific diagnostics (e.g. for use in shock experiment to allow for VISAR Barker and Hollenbach (1972) measurements on thin samples).
@@
The image is recorded on the in-air component of the MXI, an X-ray camera that is placed behind the MEC chamber. The camera can be placed between 1.3 m to 4.3 m behind the sample plane. For the large propagation distance, a fly tube is inserted and terminated by a 125 m thick kapton window to maintain low scattering and absorption transport conditions for the X-ray beam. In general, the farthest distance is preferred to have the largest magnification, although some application may benefit from the shorter distance (i.e., ptychography Schropp et al. (2010)). For most experiments, we have used an indirect X-ray imaging Optique Peter microscope Optique Peter , consisting of three options of Ce:GAG scintillators and three options of high-numerical aperture infinity-corrected optical microscope objectives (2, 4 and 10 magnification) to further chose between high-magnification or large field-of-view. This assembly is coupled to an Andor Zyla 5.5 sCMOS optical camera, which has a pixel size of 6.5 m. Direct imaging cameras (e.g., ePix or Jungfrau) can be used instead, however the resolution is limited by the pixel size (100 m for ePix10k, 50 m for ePix100 and 75 m for Jungfrau). In addition, multiframe X-ray cameras developed by hCMOS Advanced hCMOS Systems with pixel size of 25 m, Hart et al. (2019); Looker et al. (2020); Hodge et al. (2022) and capable of capturing multiple image frames spaced as little as 2 ns apart, can also be used, in conjunction with the LCLS multi-bunch operation mode Decker et al. (2022); Halavanau et al. (2021). An advantage of direct detection compared to scintillator based detection is that these cameras are orders of magnitude more sensitive, and therefore can be used with thicker, higher Z, more absorbent samples.
@@ III Phase Contrast Imaging geometry
The MXI, with a lens set of focal length , is placed in the MEC vacuum chamber and focuses the LCLS beam to approximately 100 nm. A certain distance, , behind the focus, a sample is placed in the divergent beam. The X-ray beam propagates to a detector at a distance behind the sample, leading to a geometric magnification of the sample of . Due to the long propagation distance of the X-ray, phase difference induced by the sample can readily be seen in the image (e.g., clear fringes appear at sharp phase boundaries). Iterative algorithms can be used to retrieve the phase, and if this is successful, the full complex transmission function (i.e., absorption and phase shift) can be calculated. Experiment in this geometries have been fielded at MEC, yielding scientific results on shockwave propagation in diamond Schropp et al. (2015) and the kinetics of phase transformations in silicon Brown et al. (2019). The setup can be easily combined with X-ray diffraction. Moving the lenses of the MXI closer to the target reduced the field of view of the image, and therefore the area where X-rays hit the target. In this way, diffraction from an area of only a few microns in size can be obtained, and different crystallographic phases in the sample can be identified Brown et al. (2019). A typical image of a laser-generated shock wave traveling through a sample can be seen in Fig. 4 and more examples have been published in e.g., Schropp et al. (2015); Nagler et al. (2015, 2016)
@@ IV Direct Imaging geometry
While the PCI geometry described above has great potential when a full phase retrieval can be performed, it proves to be challenging for HED samples. Without such a phase retrieval, the resolution can be severely limited, and an alternative setup that directly images a sample on the detector has been developed and used at MEC. A sketch of the setup is shown in Figure 5.
@@
In this geometry the Be lenses are placed behind the sample such that the distance to the sample, , and the distance to the camera, , matches the well known thin lens formula , with the focal length of the Be lenses. Such a setup allows for imaging intensity contrast of the sample with sub-micron resolution, even when phase reconstruction is not possible.
@@
The MXI instrument was tested in the direct imaging geometry at 8.2 keV, 9.5 keV and 18 keV photon energy. For 8.2 keV and 9.5 keV, a set of 25 Beryllium lenses with a curvature of 50 m were used. At 8.2 keV (9.5 keV), the set has a focal length of 204 mm (264 mm) and was placed 214 mm (281 mm) downstream of the target, leading to an image at the detector, placed 4.3 m behind the lenses, with a magnification of approximately 20 (15). The position of the lenses was tuned to get the best imaging resolution, using resolution targets. Images of the resolution target are shown in Fig.6. Line widths as small as 200 nm can be clearly resolved.
@@
At photon energies as high as 18 keV, the setup and alignment is more sensitive since many more Be lenses are needed. In addition, the shortest focal length that can be used is generally larger, leading to a smaller magnification. Nevertheless, the higher photon energy allows using either thicker targets, or targets containing higher Z-elements, while maintaining enough transmission to collect high-quality images. We demonstrated its use with a lens stack of 98 lenses of 50 m radius of curvature at their apex, leading to a focal length of 266 mm and a magnification of 15. Measurement on resolution targets showed a resolution similar to 8.2 keV and 9.5 keV.
@@
Examples of the use of the MXI in direct imaging geometry are detailed below. In Figure 7 a shockwave in a silicon sample, driven by the MEC short pulse laser is depicted.
This geometry is compatible with both the long pulse and the short pulse laser systems available at MEC, as well as a suite of X-ray and optical diagnostics (e.g. X-ray Diffraction, VISAR, X-ray spectroscopy, optical polarimetry) making it an ideal technique to obtain a complete picture of the dynamics happening in HED samples.
@@
The imaging setup can also be used to measure the spot size of the standard beryllium focusing lenses that are in the MEC beamline. These lenses sit approximately 4.2 m upstream of the center of the MEC target chamber (see Nagler et al. (2015) for details), and therefore a stack with focal length between 3.9 m and 4.4 m is generally used for standard MEC experiment, with a best focus around 1.2 m. Wavefront measurement have in the past been performed of such lens stack, showing significant spherical aberrations, generating ring-patterns around the focus Seaberg et al. (2019). The MXI direct imaging geometry can be used to image the intensity profile of the X-ray beam on the sample. In Figure 8, we show the intensity profile around the focus of our beamline lenses as imaged by the MXI instrument.
The airy-like ring pattern can be clearly seen. This type of measurement is of importance for all optical pump – X-ray probe experiments where X-rays are usually being used out of focus in the hope of illuminating a more uniform area. Our high-resolution imager shows, through the measurement of the X-ray caustic profile, that it might not be the case.
Table LABEL:tab:mxi_spec summarizes the working envelop of the MXI and its expected performances in direct imaging mode. The device can provide high resolution X-ray imaging over the full X-ray energy range available at MEC, albeit at varying resolution and magnification which are a function of the incoming photon energy. The total magnification is a combination of the X-ray magnification provided by the Be CRL in vacuum (this mainly constraints the diffraction limited performances), and the optical magnification provided by the objective used in the Optique Peter assembly (this allow for oversampling of the X-ray images).
@@
V Talbot Imaging
A special capability of using the MXI in direct imaging mode is his compatibility with Talbot interferometry. In this configuration, the phase accumulated by the X-ray pulse when transporting through the sample can be measured. When coupled to a phase retrieval method, this configuration has the advantage to help the removal of artifacts introduced by the use of the Be CRL. These artifacts are introduced by the density inhomogeneities of each transmissive Be lens from the stack. Some of these structures can be treated by a simple white field correction (i.e., dividing the image of the sample by the white field of image when no sample is present) or by more sophisticated analysis method such as Principal Component Analysis Hagemann et al. (2021); Hodge et al. (2022). This is the analysis that is performed in the images in Figures 4, 6 and 7. However, a large enough number of white field images (the image of the X-rays on the detector when no sample is present) needs to be accumulated in order to improve the correction with a single shot. This is necessary because the Be CRL are chromatic and the FEL bandwidth jitters, which means that a small change of the incoming central photon energy changes the focal length of the stack, effectively changing the magnification onto the detector. A division of a single shot by a white field can thus be challenging. Talbot shearing interferometry allows to mitigate these lens artifacts and chromaticity.
Talbot interferometry has been widely used at FELs to determine the wavefront of the of the X-ray beam, and its focal properties in a single shot Seaberg et al. (2019); Liu et al. (2018); Matsuyama et al. (2012). The method uses a two dimensional phase diffraction grating placed in the beam. At a specific distance behind the grating an “image” of the grating is reproduced. Analysis of the distortions, typically using Fourier methods Takeda et al. (1982) yields the differential phase in two orthogonal dimensions. These differential phases can be integrated to yield the full wavefront of the beam, and back-propagating to the focus yields the focal properties. It has been an important technique to quantify the focus in experiments where X-ray foci of order 100 nm were required.
We have combined this Talbot phase measurement technique with the direct imaging. The setup, which can be seen in Figure 9, is essentially the same as the direct imaging setup, with the addition of a Talbot grating. A checkerboard -phase grating with a pitch of 22.5 m was placed at the first Talbot distance in front of the imaging camera.
The raw image of the a shock propagating through Silicon with this setup is shown in Figure 10 10(a). Both the elastic precursor, as well as the phase transformation to denser shocked state of Si, can be seen in the raw image. Overlayed on this image is the grid pattern introduced by the Talbot grating (an zoom region is shown in the inset) that is used to measure the phase. The two dimensional fringe pattern is analyzed by standard Fourier methods Takeda et al. (1982). Since Talbot interferometry is a shearing interferometry method, the Fourier analysis results in the differential phase in the horizontal or vertical direction depending on which Fourier peak is chosen.
There are many ways Takeda et al. (1982); Seaberg et al. (2019); Nagler et al. (2017); Makita et al. (2020) to integrate the differential phase and get the phase which each have their advantages and trade-offs. For the purpose of demonstration, in this paper, we use a fast and simple method Morgan et al. (2012). Since at the first Talbot order the shear is small, we can approximate this as:
We then use the Fourier relation: